Group Selection Page
HELP
Controls Group
Data Analysis Group
ITDM
WP76
DET
Vacuum
EEE
WP78
WP74
FXE
SCS
SPB
MID
HED
SQS
SXP
Sample Environment
Operation
PSPO
XO
TS
Migrated2Zulip
|
Technical integration and coordination
SASE1 Status and SASE3 Readiness
|
Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther |
Not logged in |
 |
|
Message ID: 327
Entry time: 23 May 2022, 09:54
|
Author: |
Frederik Wolff-Fabris |
Location: |
General |
Severity: |
Information |
Subject: |
Notes XO Integration Meeting 23.05.2022 |
|
|
General Announcement:
- PRC reports the operation during the weekend run smoothly:
Tests for crosstalk between SA1/SA3 were done on FRI 23:00-00:30 (SA3 could keep their spatial overlap);
Desy phone numbers were not working on SUN morning for about 1 hour;
Please instruments shall post at least one entry per day at the Operations-2022 e-log.
CTRL:
- Karabo framework for datalogger will be rolled out today at 13:00, taking 10-15 mins and transparent to the users.
DA:
- Nothing to report.
EEE:
- Go today at SA3 PP-laser to solve issue with timing system (EEE and LAS around 11:30AM).
- MID mono motor current was changing - under investigation by EEE.
VAC:
- Nothing to report.
XPD:
- ZZ ongoing in XTD10 for motor end switches of the Imager.
DET:
- Nothing major to report.
ITDM:
- Gateway: memory usage was too high; some old Karabo sessions had to be closed as were using too memory. That prevent JDDD to be started.
XRO:
- Can machine give beam to HED this night for beamline alignment.
LAS:
- Nothing to report.
UNSYS:
- Nothing to report.
SPB:
- Air conditioning issue with laser hutch; Power plug without electricity. Tickets were issues. XO advices SPB/SFX to call TS.
FXE:
- Not present.
MID:
- All fine at MID.
HED:
- Tomorrow start commissioning with long pulse trains (every 1000th to start around lunch morning after beamline alignment). Other instruments shall not change pulses without calling BKR.
SCS:
- Nothing to report.
SQS:
- Nothing to report. |